Cinco De Mayo has a target debt-equity ratio of 0.56. (Hint: If youre not sure what to do about debt-equity ratio, watch m
Answer #1

WACC = Weight of debt * Pretax cost of debt * (1 - Tax) + Weight of Equity * Cost of Equity

Debt-Equity Ratio = 0.56

Debt/Equity = 0.56

This implies, if Equity = 1, Debt = 0.56

Weight of debt = 0.56/(1 + 0.56) = 0.359

Weight of equity = 1/(1 + 0.56) = 0.641

WACC = 0.359 * 11% * (1 - 32%) + 0.641 * 19%

WACC = 2.69% + 12.18%

WACC = 14.86%